OpenStudy (anonymous):

Help!!!! An aquarium sells cubic fish tanks of different sizes. The length of the small size fish tank is 3 feet. The dimensions of the jumbo size fish tank are double the dimensions of the small size fish tank. Which expression can be used to find the ratio of the volume of the small size fish tank to the volume of the jumbo size fish tank?

OpenStudy (anonymous):

v / V = (3^3)/ (6^3) = 27/ 216 = 1/8 So, the ratio is 1 : 8

OpenStudy (anonymous):

Thanks. That was a big help.

OpenStudy (anonymous):

You're welcome